AFIT/???/???/??-?

... THESIS

First Author

... AFIT/???/???/??-?

Approved for public release; distribution unlimited The views expressed in this thesis are those of the author and do not reflect the official policy or position of the Department of Defense or the United States Government.

AFIT/???/???/??-?

...

THESIS

Presented to the Faculty of the ... of the Air Force Institute of Technology Air University In Partial Fulfillment of the Requirements for the Degree of Master of Science

First Author, ... ...

December 1996

Approved for public release; distribution unlimited AFIT/???/???/??-?

...

First Author, ... ...

Dr. Advisor Thesis Advisor Date Dr. Member Committee Member Date Maj. Member Committee Member Date

Preface First Author

iii

Table of Contents

Page

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . iii List of Figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . v List of Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . vi List of Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . vii List of Abbreviations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . viii Abstract . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ix I. This is the first chapter title . . . . . . . . . . . . . . . . . . . . . . . 1-1

1.1 The first section . . . . . . . . . . . . . . . . . . . . . . . . . 1-1 1.2 The second section . . . . . . . . . . . . . . . . . . . . . . . . 1-1

II. The second chapter title . . . . . . . . . . . . . . . . . . . . . . . . . . 2-1

2.1 This is a section in the second chapter . . . . . . . . . . . . . 2-1

2.1.1 A subsection . . . . . . . . . . . . . . . . . . . . . . 2-1

Appendix A. First appendix title . . . . . . . . . . . . . . . . . . . . . . . A-1

A.1 In an appendix . . . . . . . . . . . . . . . . . . . . . . . . . . A-1

Vita . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . VITA-1

List of Figures

Figure Page

List of Tables

Table Page

List of Symbols

Symbol Page

List of Abbreviations

Abbreviation Page AFIT/???/???/??-?

Abstract

Abstract goes here.

...

I. This is the first chapter title

... text ... 1.1 The first section

... text ... 1.2 The second section

... text ...

1-1

II. The second chapter title

... text ...

2.1 This is a section in the second chapter ... text...

2.1.1 A subsection. Note: subsection’s and below should be printed with some sort of punctuation. A period is automatically supplied if you don’t supply some punctuation.

... more text ...

2-1

Appendix A. First appendix title

A.1 In an appendix This is appendix section A.1. Note: I highly recommend you create each chapter in a separate file including the

\chapter command and \include the file. Then you can use \includeonly to process selected chapters and you avoid having to latex/preview/print your entire document every time.

A-1

Vita

Insert your brief biographical sketch here. Your permanent address is generated automatically.

Permanent address: {... \atop ...}

VITA-1